OpenStudy (babyslapmafro):

Please help me get started with evaluating the following integral (click to see).

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\frac{ 1 }{ xln^3(x) }dx\]

OpenStudy (babyslapmafro):

What method do I use to solve this?

OpenStudy (anonymous):

put lnx=t diff it u will get 1/x=dt/dx dt=dx/x